Forecast the demand for the rice for a country for the year 2021 on the basis of 7-year



moving average and least square method, data given in table:



Year 2014 2015 2016 2017 2018 2019 2020



Population (millions) 10 12 15 20 25 30 40



Rice consumed (million tonnes) 40 50 60 70 80 90 100

1
Expert's answer
2021-12-22T13:55:31-0500

"Solution"

Moving average

"\frac{490}{7}=70"


Least square method


"n=7,\sum\ x=152, \sum\ y=490,\sum \ xy=12000,\sum\ X^2=3994"

"b=\frac{n.\sum xy-\sum x.\sum y}{n.\sum x ^2-(\sum x)^2}"


"a=\frac{\sum y \ - \ b.\sum x}{n}"


"b=\frac{9520}{4854}=1.96"

"a=\frac{192.08}{7}=27.44"


"Y=a+bx=27.44+1.96x"



"Example \ assume \ 2021\ population =50m\\consumption \ will\ be \\

Y=27.44+(1.96\times 50=125.44"